### v2

# View-Centric Multi-Object Tracking with Homographic Matching in UAV Scene

Abstract
In this paper, we address the challenge of multi-object tracking (MOT) in moving Unmanned Aerial Vehicle (UAV) scenarios, where irregular flight trajectories, such as hovering, turning left/right, and moving up/down, lead to significantly greater complexity compared to fixed-camera MOT. Specifically, changes in the scene background not only render traditional frame-to-frame object IOU association methods ineffective but also introduce significant view shifts in the objects, which complicates tracking. To overcome these issues, we propose a novel universal HomView-MOT framework, which for the first time, harnesses the view Homography inherent in changing scenes to solve MOT challenges in moving environments, incorporating Homographic Matching and View-Centric concepts. We introduce a Fast Homography Estimation (FHE) algorithm for rapid computation of Homography matrices between video frames, enabling object View-Centric ID Learning (VCIL) and leveraging multi-view Homography to learn cross-view ID features. Concurrently, our Homographic Matching Filter (HMF) maps object bounding boxes from different frames onto a common view plane for a more realistic physical IOU association. Extensive experiments have proven that these innovations allow HomView-MOT to achieve state-of-the-art performance on prominent UAV MOT datasets VisDrone and UAVDT.
1 Introduction
Figure 1: The irregular flight trajectory of the UAV causes the captured objects to exhibit a richer variety of view changes relative to a fixed shooting scene, posing greater challenges for MOT algorithms.
Multi-object Tracking(MOT) has been a classical and longstanding task in various wide applications [29, 30]. It is aimed to address the cross-frame trajectory of each object in video frames. Various MOT paradigms have been proposed for conventional static scene captured by fixed cameras [2, 21, 40]. Recently, benefited from the advancement of Unmanned Aerial Vehicle (UAV) technologies, the accessibility and analysis of scenes under moving UAV views is opening new horizons for the community [22, 56].
Figure 2: Tracking failures caused by irregular UAV flight states (turning right, hovering) leading to the ineffectiveness of ordinary IoU association. For convenience, we directly map the object box coordinates from the previous frame to the next frame (Box Position Line for clearly illustrating this process). As seen the IoU between the box of the same object across two frames (Directly Mapped Box of the previous frame and Originally Detected Box in the next frame) is very small (may be zero), causing tracker to easily lose tracking. The proposed Homographic Matching Filter utilizes Homography spatial relationship between two frames to map the previous frame box homographically onto the next frame before performing IoU association (Homographic Mapped Box of the previous frame, and Originally Detected Box in the next frame), clearly showing a substantial increase and a more reasonable IoU (physical IoU in the real world).
Compared to the fixed scenes captured by conventional stationary cameras, due to the movement of the UAV itself in a moving UAV scenario, the captured scene is also in motion, making it more difficult to track multiple objects with different motion states within the scene. Specifically, this manifests in two main ways: (1) During the irregular trajectories in the UAV flight, the variable shooting angles of the UAV lead to a rich diversity of view changes in the captured objects, posing a significant challenge for object ID discrimination learning during the tracking process, as shown in Figure 1. (2) As the overall scene changes, it becomes challenging to capture the actual motion states of the objects, leading to numerous tracking losses. Specifically, as the UAV moves, the relative coordinates of the objects against the background become highly variable, introducing significant errors and uncertainties in the frame-to-frame ID matching process during tracking. As shown in the examples in Figure 2, we can see that the vehicles framed in two consecutive frames have only moved slightly relative to the ground in the real world. If captured by a fixed camera with an unchanging scene, these objects could be easily matched through IoU association, which is widely used in classical MOT algorithms. However, due to the UAV’s potential for various state changes such as hovering (row 2 in Figure 2), turning left/right (row 1 in Figure 2), and moving up/down, the IoU of the absolute coordinates between the two frames (calculated between the directly mapped box of previous frame and the originally detected box in next frame) can be very small (even zero), resulting in match losses.
In this paper, we envision utilizing Homography property of spatial scenes to address the aforementioned issues and propose a novel View-Centric Multi-Object Tracking with Homographic Matching (HomView-MOT) in Moving UAV scene. Firstly, for each UAV-captured video sequence, we regard it as multi-view shooting of one or more scenes. Therefore, we initially compute the Homography transformation matrix between adjacent frames. Notably, to enhance tracking efficiency, we avoid calculating between every two adjacent frames and instead perform skip-frame computation, only calculating the Homography matrix for every $h$ frames apart. We design an efficient Fast Homography Estimation (FHE) algorithm to reasonably estimate the Homography matrix for every two frames within this interval. Secondly, based on the inter-frame scene Homography matrix, we present the Homographic Matching Filter (HMF), which can map detected objects within them to the same viewpoint space to calculate the "physical IoU in the real world" of objects between two frames (as shown in Figure 2), thereby improving the precision of track matching. Finally, to address the issue of object view changes during UAV flight, we propose View-Centric ID Learning (VCIL). By combining the inter-frame scene Homography matrix, we design Homographic Slot Attention (HSA) to explicitly extract view-centric slots for each frame and correlate views between adjacent frames, thus updating object ID features to cope with view changes during UAV flight. By integrating these designs, we make HomView-MOT an efficient and effective multi-object tracker.
Essentially, HomView-MOT is a specialized framework designed for "multi-object tracking in moving scenarios". In this paper, we validate our approach on two typical UAV MOT datasets, including VisDrone 2019 [56] and UAVDT [8]. The main contributions of this paper are as follows:
  • For the first time, we address the problem of multi-object tracking (MOT) in moving scenarios by integrating the concept of Homography from different views, introducing Homographic Matching and View-Centric concepts into the field and proposing a universal HomView-MOT framework.
  • Specifically, we begin by employing a Fast Homography Estimation (FHE) algorithm to rapidly estimate the Homography matrix between any two video frames. Based on this, we perform object View-Centric ID Learning (VCIL), utilizing the Homography of multiple perspectives to learn cross-view ID features. Furthermore, leveraging the Homography matrix, we design a Homographic Matching Filter (HMF) to map the object bounding boxes from different frames onto the same view plane, allowing for more accurate physical IoU associations that closely mirror the real world.
  • Through extensive experiments, we validate that the above designs enable HomView-MOT to achieve state-of-the-art performances on two typical UAV MOT datasets, VisDrone and UAVDT.
2 Related Work
Multi-Object Tracking. Classical MOT algorithms [2, 34, 4, 9] generally adopt a tracking by detection framework, where the first step is to detect all object boxes in each frame, and the second step involves object ID matching between adjacent frames through data association[42, 34]. In the first detection step, related algorithms such as SORT[2], Deep SORT[34], and others typically use classical detection models like Fast-RCNN[10] and YOLO[26]. Benefiting from the recent progress in deep learning[60, 58, 57, 59, 51, 54, 52, 53, 55, 18, 16, 17, 15, 13, 12, 14, 49, 46, 48, 47, 38, 50, 37, 39, 31, 32], recent research has also introduced vision transformer detectors into MOT to achieve higher performances, as seen in TransTrack[28], TrackFormer[24], and similar approaches. In the second step of data association, traditional MOT algorithms usually rely on Matching Algorithms based on IoU association between target boxes in adjacent frames. This method is quite effective in static scenes captured by fixed cameras. However, in UAV footage and the like, the moving scene background, causing different frames to be on different viewing planes, can render the conventional IoU association method inaccurate. To address this, UAVMOT[22] proposes an Adaptive Motion Filter to judge the different motion states of the UAV and process them differently. In this paper, we propose a Homography Matching Filter that, based on Homography transformation, maps the object boxes from adjacent frames in different perspectives into the same viewpoint space before performing IoU association to enhance the robustness of tracking.
ID Embedding Assistance. To enhance robustness, many MOT frameworks propose adding an embedding based on ReID learning on top of detection, which is used to improve multi-object data association[43, 22]. In scenes captured by stationary cameras, targets generally present a very stable viewpoint, making ID learning comparatively simple and straightforward. However, in the case of UAVs, due to the variability of shooting angles, an object may exhibit a rich array of viewpoints. Therefore, we further propose to explicitly learn the view characteristics of targets to optimize the feature representation of object IDs.
3 Method
3.1 Overview
Our proposed HomView-MOT, as shown in Figure 3, follows the classic tracking-by-detection paradigm. Given a video sequence acquired by a moving UAV, the goal of MOT is to obtain the categories, bounding boxes, and tracking IDs of objects in the video. First of all, we present a Fast Homography Estimation (FHE) algorithm to calculate the Homography transformation matrix for any adjacent frames, which is used for scene view projection. Taking the adjacent $t$-th and $t+1$-th frames as an example, these two frames complete the detection of objects through the detection head of a shared feature extraction network and output bounding boxes (specifically, obtained by the heatmap and the width and height of the box), as well as objects ID embedding features. Next, we propose a View-Centric ID Learning (VCIL) module to explicitly learn the multi-view features of each object and update the object ID features. Then, in the tracking phase, we propose the Homography Matching Filter (HMF), which maps the object boxes from adjacent frames onto the same view planar space for IoU association, combined with ID feature similarity, and finally sends them to the Matching Algorithm to output the final tracking results.
Figure 3: The overview of our proposed HomView-MOT, adhering to the classic tracking-by-detection paradigm. It begins with a Fast Homography Estimation (FHE) algorithm that rapidly calculates the Homography transformation matrix between any two adjacent frames for view scene projection. With a shared feature extraction network, we obtain the objects detection boxes ($B_{t}$, $B_{t+1}$) and ID features ($ID_{t}$, $ID_{t+1}$). Subsequently, the View-Centric ID Learning (VCIL) is utilized to explicitly assimilate the multi-view information and update the ID features. For tracking, the Homography Matching Filter (HMF) is employed to project the object boxes from adjacent frames onto a common view planar space for IoU association. This is coupled with ID association and processed through Matching Algorithm to produce final tracking results.
3.2 Fast Homography Estimation for View Projection
As described in Sec. 1, when the UAV flight state changes (e.g. hovering, moving up/down, turning left/right) occur, the captured scene background will undergo significant changes, making adjacent frames appear as different view planes over the same scene. To facilitate cross-view learning and IoU association for object IDs, we propose to perform view projection for these two planes, in the form of Homography matrix which is extensively applied to various 3D applications [44]. Generally, the Homography matrix for every two adjacent frames can be calculated with the Homography Estimation (HE) algorithm. For instance, as shown in Figure 4, we use a shared SuperPoint [6] network to extract keypoints from both frames separately. The distribution of these keypoints is able to represent the view information of the scene [41, 20, 45, 19]. Then, we can estimate the Homography matrix based on the matching of keypoints between the two frames by the SuperGlue algorithm [27].
Figure 4: One example of the calculation of Homography matrix. The lines indicate the matching of keypoints between the two frames.
Figure 5: Illustration of the sampled key frames and non-sampled frames in the video sequence, where $h=5$, for example.
However, directly performing Homography estimation calculation for every two frames is very time-consuming and would greatly reduce tracking efficiency. Therefore, we propose a Fast Homography Estimation (FHE) method, where we only need to directly calculate the Homography matrices for a limited number of key frames, while the Homography matrices for other adjacent frames can be fast estimated using the ones of the key frames. Specifically, we utilize a uniform sampling process to select the key frames from the input video sequence $I_{t}(t\in[1,T])$, the sampling interval is set to $h$. We only perform a direct calculation of the Homography matrices $\{H_{t,t+h},H_{t+h,t+2h},...\}$ between the sampled frames, where $H_{t,t+h}$ indicates the Homography matrix from the $t+h$-th to $t$-th frame. For any two frames , based on the Derivation 1 in the appendix, we prove that their Homography matrix can be reasonably estimated through the derivation of Homography matrix of their adjacent key frames.
Specifically, as shown in Figure 5, We select three highly representative frames $I_{t_{1}},I_{t_{2}},I_{t_{3}}$ to illustrate the computation process of the Homography matrix between any two frames, where $k_{1}<t_{1}<t_{2}<k_{2}<t_{3}<k_{3}$, and $k_{1},k_{2},k_{3}$ are the sampled frames. The core calculations are as follows:
  1. 1.
    For the same frame, that is, for any given frame $t$, according to the property of Homography, we have $H_{t,t}=E$, where $E$ is the identity matrix.
  2. 2.
    Between any two sampled key frames, we compute the Homography matrix $H_{k_{2},k_{1}},H_{k_{3},k_{1}},H_{k_{3},k_{2}}$ using the naive Homography Estimation algorithm.
  3. 3.
    For any two frames $t_{1},t_{2}$ within the same sampling interval:
    • From a non-sampled frame to the previous sampled frame: $H_{t_{1},k_{1}}=\alpha_{t_{1},k_{1}}H_{k_{2},k_{1}}$ and $H_{t_{2},k_{1}}=\alpha_{t_{2},k_{1}}H_{k_{2},k_{1}}$, where $\alpha$ is scaling factors based on the proportion of the overall displacement of key points between the two frames to the overall displacement of key points between the adjacent key frames $k_{1},k_{2}$. More details refer to the appendix.
    • From a sampled frame to a non-sampled frame: $H_{k_{2},t_{1}}=H_{k_{2},k_{1}}H^{-1}_{t_{1},k_{1}}$ and $H_{k_{2},t_{2}}=H_{k_{2},k_{1}}H^{-1}_{t_{2},k_{1}}$.
    • From a non-sampled frame to another non-sampled frame: $H_{t_{2},t_{1}}=H_{t_{2},k_{1}}H_{t_{1},k_{1}}^{-1}$.
  4. 4.
    For any two frames $t_{1},t_{3}$ in different sampling intervals:
    • From a non-sampled frame to any frame in the previous sampling interval: $H_{t_{2},t_{1}}=H_{t_{2},k_{2}}H_{k_{2},t_{1}}$ and $H_{t_{2},k_{1}}=H_{t_{2},k_{2}}H_{k_{2},k_{1}}$. The calculation of $H_{t_{2},k_{2}}$ is similar to the previous criterion.
The aforementioned computation process involves only interpolation calculations and matrix multiplication, which ensures high computational efficiency in practical implementation. Through the derivations described above, we can rapidly and cost-effectively calculate the Homography matrix between any two frames.
3.3 View-Centric ID Learning
Building upon the previously calculated Homography matrix between adjacent frames, we can facilitate multi-view learning of object IDs.
In MOT, ID learning is critical for identifying object ID information to facilitate data association. Due to the dynamic nature of UAV-MOT scenes where both the background and objects may be moving, the objects may exhibit a rich variety of view changes. Therefore, we propose a novel View-Centric ID Learning approach, which is inspired by the Object-Centric Learning framework [23]. Specifically, in combination with the Homography of the view plane in different frames, we first propose Homography Slot Attention, which uses "slots" [23, 35] to explicitly learn the view information of an object, referred to as View-Centric Slots. We correlate the extracted view features of objects from adjacent frames to obtain updated view slots, and then through a decoding process, we reconstruct a new ID embedding and update the object’s ID embedding feature.
Figure 6: The illustration of View-Centric ID Learning. The $\widetilde{S_{t}}$ and $\widetilde{S}_{t+1}$ indicate the randomly initialized view-centric slots.
Within the foundational concept of Slot Attention, The Slot Attention [23, 35] maps from a set of input feature vectors to a set of output vectors, referred to as slots. Each vector in this output set can describe an object or an entity in the input. Building on this concept, we propose Homographic Slot Attention (HSA), which leverages the view-based Homography matrix and iterative attention to map its inputs to view-centric slots. At each iteration, slots compete to describe views of the input through a softmax-based attention mechanism and Homography matrix. As shown in Figure 6, given two frames $t,t+1$, we describe their one iteration of HSA on the ID feature extracted by the detection head as,
${\rm HSA}_{t}$ $={\rm Softmax}(\frac{1}{\sqrt{C}}k(F_{t}^{ID})\times(H_{t,t+1}\cdot q(S_{t})^{\top}))\times v(F_{t}^{ID})$ (1)
${\rm HSA}_{t+1}$ $={\rm Softmax}(\frac{1}{\sqrt{C}}k(F_{t+1}^{ID})\times(H_{t+1,t}\cdot q(S_{t+1})^{\top}))\times v(F_{t+1}^{ID})$
where $q(\cdot),k(\cdot),v(\cdot)$ are the linear projections. $F_{t}^{ID},F_{t+1}^{ID}$ are the object ID features in frame $t,t+1$, respectively. $S_{t},S_{t+1}$ are the View-Centric Slots in frame $t,t+1$, respectively. $\top$ is the matrix transposition operator. $C$ is the feature dimension. Then we correlate the cross-frame views by,
$S^{\prime}={\rm Softmax}(\frac{S_{t}\times S_{t+1}^{\top}}{\sqrt{C}})\times S_{t+1}$ (2)
then a MLP-decoder $\mathcal{D}$ is used to reconstruct visual feature ${F^{ID}}^{\prime}$ from the correlated slots $S^{\prime}$. Finally, ${F^{ID}}^{\prime}$ is correlated with $F_{t_{2}}^{ID}$(current frame) for ID feature updating, in the cross-attention manner.
3.4 Homography Matching Filter
Leveraging the Homography matrix, we design a Homographic Matching Filter (HMF) to map the object bounding boxes from different frames onto the same view plane, allowing for more accurate physical IOU associations that closely mirror the real world.
In UAV video sequences, the movement of objects is not limited to linear motion but becomes nonlinear and irregular due to the coupling of the UAV’s and the object’s own motions. The traditional Kalman filter [33] struggles with this irregular motion, so we introduce the HMF to handle the complex movements associated with UAVs. The HMF module adaptively projects cross-frame boxes onto the same view plane, making IoU association more regular and normalized, which allows for accurately completing the object ID association, relative to the real world.
Figure 7: The illustration of Homography Matching Filter (HMF). We apply Homographic mapping between two adjacent frames and compute the Homographic IoU for each, then take the average of the IoUs to enhance robustness.
As shown in Figure 7, to ensure robustness, we perform cross-projection between the object bounding boxes $\{B_{t}\}$ and $\{B_{t+1}\}$ of two adjacent frames $t,t+1$. For each pair of $B_{t}$ and $B_{t+1}$, the calculation of HMF is,
${\rm HMF}(B_{t},B_{t+1})=({\rm IoU}(H_{t,t+1}B_{t},B_{t+1})+{\rm IoU}(H_{t+1,t}B_{t+1},B_{t}))/2$ (3)
Comparison with existing Motion Filters. Previous motion filters in UAV MOT, such as Adaptive Motion Filter (AMF) in [22], typically rely on the spatial relationships of objects to their surroundings to determine the UAV’s movement state. However, this approach is heavily dependent on manual design and lacks robustness when the UAV’s motion state is complex, as well as obviously affect the tracking efficiency. In contrast, our proposed HMF is more concise, lightweight, and robust.
3.5 Loss Design and Overall Tracking Algorithm
Along with the basic detection loss, we also add the reconstruction loss, for reconstructing the view-centric slots to visual feature. So the overall loss is formulated as,
$\mathcal{L}=\mathcal{L}_{{\rm basic}}+\lambda\mathcal{L}_{rec}=\mathcal{L}_{{\rm basic}}+\lambda\sum_{t=1}^{T}\mathcal{L}_{CE}(F^{ID}_{t},\mathcal{D}(S_{t})),$ (4)
where $\lambda$ is loss weight. We use cross-entropy loss both in $\mathcal{L}_{rec}$.
The proposed tracking algorithm can be summarized as Algorithm 1.
Input: An UAV vide sequence $\{I_{t}\in\mathbb{R}^{W\times H\times 3}\}^{T}_{t=1}$
Output: The tracked objects $T_{t}=\{B_{t},C_{t},ID_{t}\}$
   while $t<T$ do
     if $t$ is divisible by $h$ then
       Calculated $H_{t,t+h}$ of $I_{t},I_{t+h}$ directly by Homography Estimation.
     end if
     Derive $H_{t,t+1}$ from $H_{t,t+h}$ with Fast Homography Estimation algorithm.
     Input two adjacent frames $I_{t},I_{t+1}$.
     Finish object detection, obtain the detected boxes $\{B_{t},C_{t}\}$, $\{B_{t+1},C_{t+1}\}$.
     Calculate IoU association with Homography Matching Filter.
     Obtain ID embedding features $F^{ID}_{t},F^{ID}_{t+1}$, calculate the feature association.
     Update the ID feature $F^{ID}_{t+1}$ with View-Centric ID Learning.
     Ensemble IoU and feature association, perform Matching Algorithm.
   end while
Algorithm 1 HomView-MOT algorithm
4 Experiments
4.1 Datasets
VisDrone2019: The VisDrone2019 dataset [56] is applied for tracking and detection from UAV perspectives. For the MOT task, the VisDrone2019 dataset comprises a training set (56 sequences), a validation set (7 sequences), and a test set (33 sequences, with test-challenge containing 16 sequences and test-dev comprising 17 sequences). In each frame, every object is labeled with a bounding box, category, and tracking ID. The VisDrone2019 dataset encompasses ten categories: pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, and tricycle. For the multi-object tracking evaluation, only five object categories are considered, namely car, bus, truck, pedestrian, and van.
[…]
MOTDT [5] - -0.8 21.6 87 1196 44548 185453 1437 3609
SORT [2] 23.5 14.0 38.0 506 545 80845 112954 3629 4838
IoUT [3] 27.3 28.1 38.9 467 670 36158 126549 2393 3829
GOG [11] 2.0 28.7 36.4 346 836 17706 144657 1387 2237
MOTR [40] 7.5 22.8 41.4 272 825 28407 147937 959 3980
TrackFormer [24] 7.4 25 30.5 385 770 25856 141526 4840 4855
UAVMOT [22] 12.0 36.1 51.0 520 574 27983 115925 2775 7396
ByteTrack [42] 11.4 52.3 68.3 769 325 24156 110342 2230 4001
U2MOT [21] 19.4 52.3 69.0 - - - - 1052 -
HomView-MOT 20.8 54.2 75.1 870 224 31461 72600 1073 2832
Table 1: Comparison with state-of-the-arts on VisDrone test-dev dataset.
[…]
CEM [25] -6.8 10.1 94 1062 64373 298090 1530 2835
SMOT [7] 33.9 45.0 524 367 57112 166582 1752 9577
GOG [11] 35.7 0.3 627 374 62929 153336 3104 5130
IOUT [3] 36.6 23.7 534 357 42245 163881 9938 10463
CMOT [1] 36.9 57.5 664 351 69109 144760 $1$ 3656
SORT [2] 39 43.7 484 400 33037 172628 2350 5787
DSORT [34] 40.7 58.2 595 338 44868 155290 2061 6432
MDP [36] 43.0 61.5 647 324 46151 147735 541 4299
UAVMOT [22] 46.4 67.3 624 221 66352 115940 456 5590
HomView-MOT 48.1 70.7 694 234 58834 117691 427 6003
Table 2: Comparison with state-of-the-arts on UAVDT test dataset.
UAVDT: The UAVDT dataset [8] is specifically utilized for vehicle object detection and tracking, featuring three categories: car, truck, and bus. For the MOT task, it is divided into a training set (30 sequences) and a test set (20 sequences). It exclusively focuses on a single category, car. The video frames have a resolution of 1080 × 540 pixels and encompass various common scenes, such as squares, arterial streets, and toll stations.
4.2 Implementation Details
To ensure a fair comparison, we follow the experimental setup of existing works such as UAVMOT [22]. Regarding training details, the total number of epochs is set to 30, with the learning rate decaying by a factor of 10 at 10 epochs and 20 epochs, respectively. The initial learning rate is set to 7e-5. For data augmentation, we utilize random cropping and random scaling (between 0.6 to 1.3). The experiments are conducted on two GeForce RTX GPUs with a batch size of 4. In the multiple loss functions, L1 loss is used to supervise the object’s width and height. Cross-entropy loss and triplet loss are employed for handling the object ID. Additionally, we utilize the standard cross-entropy loss to supervise the object heatmaps and view slot reconstruction. The loss weight $\lambda$ is set to 0.4.
[…]
Baseline 2079 40.6
Baseline+IDFU[22] 937 43.8
Baseline+VCIL 790 51.9
Table 3: The ablation study of View-Centric ID Learning (VCIL) on VisDrone validation set.
[…]
Baseline 2079 40.6
Baseline+AMF[22] 958 44.1
Baseline+HMF 816 53.7
Table 4: The ablation study of Homographic Matching Filter (HMF) on VisDrone validation set.
Figure 8: The robustness analysis of Fast Homography Estimation (FHE). Horizontal axis: Sampling interval $h$ in the FHE algorithm. Vertical axis: Tracking accuracy (IDF1) and speed (FPS) of HomView-MOT under different settings of $h$. When $h=1$, FHE degenerates into the ordinary HE algorithm, yielding high accuracy but at a slow speed. As $h$ increases, the tracker maintains relatively high accuracy while the speed significantly improves, confirming the value of the FHE approach. However, if $h$ is too large, the calculation of the Homography matrix becomes imprecise, which affects performance.
Figure 9: The case studies of HomView-MOT. We show three UAV flight states: turning up, Hovering in the sky, and turning left. As seen that the proposed Homographic Matching Filter demonstrate better and more reasonable IoU association (relative to the real world) then traditional ordinary IoU association, by comparing the Directly Detected Boxes and Homographic Mapped Boxes.
4.3 Comparison with State-of-the-arts
We evaluate from the two most critical dimensions: accuracy and speed, on the VisDrone and UAVDT datasets.
VisDrone Following previous works[22, 21], we benchmark our method against previous approaches on the VisDrone2019 dataset for the MOT task. We train our model using both the training and validation sets and assess our approach on the VisDrone2019 test-dev set utilizing the official VisDrone MOT toolkit. As depicted in Table 1, our method attains a 54.2% score on MOTA and 75.1% on IDF1, surpassing existing techniques on the VisDrone2019 test-dev set.
UAVDT In addition to our evaluation on the VisDrone2019 dataset, we also benchmark our method against other approaches on the UAVDT test set for the MOT task. We train the proposed network on the UAVDT training set and evaluate our approach on the UAVDT test set. We present a range of metrics, including MOTA and IDF1, to compare the performance of our method with that of others. As shown in Table 2, our method registers a 48.1% score on MOTA and 70.7% on IDF1, achieving significantly enhanced results compared to existing methods.
4.4 Ablation Study
4.4.1 Effectiveness of View-Centric ID Learning
View-Centric ID Learning (VCIL) enables ID embedding to learn multi-view information, effectively enhancing the reliability of ID feature association and thereby bolstering the robustness of MOT algorithms in moving UAV scenarios. To verify the effectiveness of VCIL, we study its benefits on ID feature association using two metrics: IDs, IDF1. For fair comparison with previous methods, we follow the same baseline with [22]. Specifically, we evaluate the performance of the baseline, the baseline with the IDFU module introduced in [22], and the baseline with VCIL. The experimental results, as shown in Table 4.2, indicate that the inclusion of VCIL leads to a significant improvement across all metrics compared to both the baseline and IDFU. These results demonstrate that VCIL has a substantial positive impact on the ID feature association process, allowing the model to quickly acquire multi-view information of targets during tracking.
4.4.2 Effectiveness of Homography Matching Filter
To address the challenges posed by the irregular motion of targets in moving UAV scenarios, we present the Homography Matching Filter (HMF). To validate its effectiveness, we compare its performance with that of ordinary IoU association, the adaptive motion filter (AMF) proposed in [22], as well as against HMF itself. For fair comparison with previous methods, we follow the same baseline with [22]. Particularly, our HMF is a lightweight yet effective algorithm. The experimental results, as presented in Table 4.2, are based on two metrics: IDs, IDF1. It is evident that HMF demonstrates improved accuracy over both the baseline and AMF. Notably, in contrast to the manually intensive and complex "UAV motion mode selective + local relation filter" used by AMF, HMF is more lightweight and, while maintaining higher accuracy, also offers faster speed.
4.4.3 Robustness Analysis of Fast Homography Estimation
Within the design of HomView-MOT, the calculation of inter-frame Scene Projection through Fast Homography Estimation (FHE) is fundamental and critical. In our design, to ensure the overall efficiency of tracking, we compute the inter-frame Homography Matrix rapidly by using interval sampling coupled with FHE, based on the assumption that the scene can be reasonably considered to be moving linearly within a small sampling interval. This design ensures the efficiency and precision of our HomView-MOT. Generally, a larger sampling interval tends to increase tracking efficiency, but an overly large interval may weaken the assumption of "approximate linear motion within the interval," thus compromising tracking accuracy. Figure 8 displays the performance of tracking under different settings of $h$. It is observed that a smaller $h$ yields the highest accuracy but also reduces speed. As $h$ increases, we maintain a high level of accuracy while also achieving a higher speed, but a further increase in $h$ will lead to a noticeable decrease in accuracy. These experimental findings strongly support the aforementioned arguments.
4.4.4 Case Study
To better illustrate the effectiveness of the Homography Matching Filter, we visually present in Figure 9 the performance of HMF when the UAV-captured scene moves, causing the target trajectories to become irregular. Figure 9 depict three different UAV status changes: hovering, turning up, and turning left, showing comparison between two consecutive frames. It is evident that the objects have not moved significantly in the real world (as can be seen with reference to ground features such as lane lines), and when directly mapped, the IoU between the detection boxes of consecutive frames is very low, leading to potential match losses. However, with the mapping based on Homography, the detection boxes from frame to frame are closer together, yielding a higher IoU and thus achieving a more reliable IoU association.
5 Conclusion
Our work presents a pioneering HomView-MOT framework that successfully addresses the inherent challenges of multi-object tracking in moving UAV scenarios. By innovatively employing the principles of Homography and View-Centric learning, our approach effectively navigates the complexities introduced by UAV flight dynamics, such as irregular motion patterns and varying view perspectives. The Fast Homography Estimation (FHE) algorithm computes homography matrices efficiently, while our Homographic Matching Filter (HMF) and View-Centric ID Learning (VCIL) techniques facilitate accurate and robust tracking across changing scenes. The effectiveness of our framework is underscored by its state-of-the-art performance on the VisDrone and UAVDT datasets.
Limitation and Outlook. In fact, HomView-MOT is a universal framework designed for "multi-object tracking in moving scenarios" and is not limited to UAVs alone; in theory, it can be applied to a variety of industrial applications including different types of aircraft, handheld recording devices, underwater diving cameras, and more. Due to the current limitations of available open-source datasets, we have only been able to validate our framework in moving UAV scenarios. In the future, we look forward to gathering more datasets relevant to diverse scenarios in order to more broadly ascertain the efficacy and versatility of our method.